What is the measure of Angle C?
Angle A: 5x+12
Angle B: 6x+30
Angle C: 2x+8

Answer by mananth(16949)


5x+12 + 6x+30 + 2x+8=180 ( sum of angles of triangle)
13x+50=180
13x=130
x=10
C= 2x+8
plug x=13
13*10+8
=138
C= 138